A solution of 0.15M NH3 is only 1.1% ionized in solution. Calculate the pH. The Kb of ammonia is 1.8*10^-5


1
Expert's answer
2021-02-15T03:23:51-0500

The reaction is:-

"NH_{{3}_{(aq)}}+H_2O_{(l)}\rightarrow NH^+_{4_{(aq})}+OH^-"

Given concentration of "NH_3=0.15M"


As It ionize 1.1% only So Concentration of "NH_4^+" will be-

"[NH_4^+]=\dfrac{1.1\times 0.15}{100}=1.65\times 10^{-3}M"


Also, "k_b=\dfrac{[NH_4^+][OH^-]}{[NH_3]}"


"1.8\times 10^{-5}=\dfrac{1.65\times 10^{-3}[OH^-]}{.15}"


"\Rightarrow [OH^-]=\dfrac{0.27\times 10^{-5}}{1.65\times 10^{-3}}=1.63\times 10^{-3}M"


"P_{OH}=-log([OH^-])=-log(1.63\times 10^{-3})"

"=2.79"


As we know, "P_H+P_{OH}=14\\"

"P_H=14-P_{OH}=14-2.79=11.21"


Hence "P_H" is 11.21.